In a multi-bank system, after customers deposited additional $720 billion in the bank, M1 money supply increased by $2,160$⁢2,160 billion. Calculate the value of loans issued by the bank during the first round of loans. Assume that all the excess reserves are loaned out. Enter your answer in billions in the box below.
Answer #1

Answer -

Money multiplier = 2160/720

= 3

Money multiplier = 1/ reserve ratio

3 = 1 / reserve ratio

Reserve ratio = 1/3

Deposits = $ 720 billion

Required reserve = 720*1/3

= $ 240 billion

Excess reserves = 720-240

= $ 480 billion

Since all of the excess reserves were loaned out , loan issued in first round

= $ 480 billion